OpenStudy (anonymous):

For which values of b are the vectors (3, b), (6, b-1) linearly independent?

OpenStudy (amistre64):

terminology ... what does it mean to be linearly independant?

OpenStudy (anonymous):

Sorry,, I went on working some other probs where I didn't have to figure out the value of b. I still have quite a few though that are the same type as my posted question...explanation would be helpful. See the attachment regarding terminology.

OpenStudy (amistre64):

according to the text there; the number of vectors in the space is important; <x,y> would suggest R^2 to me. since we have 2 vectors lets set them up in the matrix by rows: \[\left( \begin{array}c 3&b\\ 6&b-1\\ \end{array} \right)\] a 2x2 matrix can easily be solved for the "det". 3(b-1) - 6(b) not= 0 id say determine the values for b and exclude them 3b -3 -6b = 0 -3b -3 = 0 -3b = 3 b = -1; this would be a bad case according to the text you provided.

OpenStudy (anonymous):

gotcha! Form a matrix with the vector, find the determinant, then determine whether or not we want that value included or excluded depending on if we're talking about independent or dependent linearity. Thanks a bunch!